OpenStudy (anonymous):

Choose one of the factors of 1331x^3 – 8y^3 Answer 121x^2 + 44xy + 4y^2 121x^2 – 22xy + 4y^2 4 11x – 2y

OpenStudy (mathmate):

Knowing that 1331x^3 – 8y^3 = (11x)^3-(2y)^3, and from A^3-B^3=(A-B)(A^2+AB+B^2) you can eliminate the second and third. Check the first one just to be sure.